Shift Summary: Owl 11/20/99

Jason LaDue (jladue@cuhep.Colorado.EDU)
Sat, 20 Nov 1999 06:04:58 -0700 (MST)


Shift Summary for owl shift, 11/20/99

On Shift: U. Erkarslan, J. LaDue

Hours of beam available at all intensities: 7.00
Hours of beam available at nominal intensity: 7.00
Hours of E799 data (special runs not included): 5.75
Hours of special runs (laser scans, muon runs...): 0.25

Accounting of lost time:
Stopping/starting runs, KFRENDS.

========================================================================
Runs ended this shift:
15017 - 3.7M - BEAM_799
15018 - 4.8M - BEAM_799
15019 - - BEAM_MUON_4E12

Tapes brought to FCC this shift:

========================================================================
Accesses during shift:

========================================================================
Overview of shift:

There is a problem with DC4X_H_R, it was reading at about~3/4 the
rate of DC4X_H_L, this lead to a redefinition of the Hyperon triggers
removing the DC4X_H_L and DC4X_H_R requirements. Some trouble with getting
a run started, but otherwise pretty smooth. Labeled UPH031-060 split
tapes.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15017 99 00:00 DC4X_H_R rate is screwy again (see previous shifts
comments).
15017 00:15 Run Stopped so that we can update the trigger file.
It removes the requirement of 4X from the Hyperon
triggers.
15018 27 00:54 After some shenanigans, finally get the correct
trigger file loaded.
15018 02:47 Tapes full, run ends. Change tapes.
03:00 UPM033 split tape mount request.
15019 03:08 Start Muon run.
15020 03:35 Start data taking after 6 KFRENDS and two laser
repeater resets.
15020 03:53 Triggers stop with a bunch of fruit busies and
VVH DYC busy. KFREND and continue.
15020 54 04:40 Pipeline errors. Init Pipeline and KFREND.

========================================================================
Follow up Notes on Previous Shift
Got new trigger files installed for Triggers 10, 11, 12.
Labeled UPH split tapes.

========================================================================
Items to Follow up on Next Shift